Molybdenum forms body-centered cubic crystals and at 20 oC the density is 10.3 g/mL. Calculate the distance between the centers of the nearest molybdenum atoms.


1
Expert's answer
2021-08-20T01:51:10-0400

Solution


Density = Z "\times" "\frac {M}{N}" "\times{l^3}"


Z is effective number of atoms for Molybdenum (Mo). = 2 .


M is Molar mass of Mo = 95.95 g/ mol.


N = 6.02 × 1023.


L is edgelength in Cm .


d = 10.3 g / "cm^3."


M = 95.95 g/ mol.


N = "6.02 \u00d7 10^{23}"


Make L subject of the formula


L = 3.16 × 10-8 Cm.